OpenStudy (anonymous):

how many molecules are in 23 moles of oxygen??

OpenStudy (anonymous):

23 moles Oxygen × 6.02 × 1023 atoms = 1.3846 × 1025 atoms Oxygen 1 mole So there are 13846000000000000000000000 atoms in 23 moles of Oxygen

OpenStudy (anonymous):

where did you got the 6.02 and 1023 out??

OpenStudy (anonymous):

No of Molecules = no of moles * Na

OpenStudy (anonymous):

Na = Avagadro number \[6.022 * 10^{23}\]

OpenStudy (anonymous):

\[6.022*10^{23} * 23 =138.506 *10^{23}\]

OpenStudy (anonymous):

ouu okk.. i got it know thankks!

OpenStudy (anonymous):

Welcome....).....Dont Forget to Give medal by clicking best answer option